Problem 1 (10pts) Jims utility function is U (x, y) = xy. Jerrys utility function is U (x,y) = 1,000xy +2,000. Tammys util
Answer #1

Answer: 1

(a) MRS is the rate at which a consumer is ready to give up some units of good 1 to get an additional unit of good 2 while maJims W (ny) = ny MUX MUY Jemys u(,y)= 10002y + 2000 Mい。 Mux - 10001 - MUу 1000x Tammys ulay = xylion) 29 - 1y? MUx MUY y-2Marjoes ulny) = (y +1000) = xy + 10000 MU 1000 MUу Pats ( Up, y)) = o.sxy - 10,000 Billys (u(ny) - 14y. I Musly MUY fran eb) Jerry, pat, Dave and Bill have the same in difference Curves as a Jim, but Tammy and Francis have different preferences fr